# The early universe is ACT-ing warm

### v1

@@ Abstract
The recently released data from the Atacama Cosmology Telescope (ACT) confirms that the primordial scalar spectrum is extremely flat. This, together with current upper bounds on the tensor-to-scalar ratio, implies that the simplest models of inflation coming from particle physics (for instance, a minimally-coupled scalar with monomial potentials) need additional ingredients in order to make them compatible with observations. Instead of invoking arbitrary new couplings or new interactions that are not protected symmetries, we argue that dissipation of the inflaton field with the radiation bath should be added as a new physical principle. Accordingly, we show that warm inflation provides the correct paradigm to explain the current observations, given very natural choices of dissipative terms.
I Spoiled for choice, yet not a single ‘principled’ one
@@
The main prototype for inflation is a scalar field slowly rolling down an almost flat potential, such that the potential energy of the field dominates over its kinetic energy, and sources the Friedmann equation. The Hubble radius remains approximately constant during this phase, leading to an increase of the size of the universe by a factor of $e^{3N}$ if inflation lasted for $N$ $e$-folds. This results in a supercooled universe which is practically empty, and one needs a phase of reheating/preheating where the energy density of the inflaton is released by creating a shower of particles in radiation form.
The above story is simple enough that theorists have successfully built a zoo of potentials ($\sim\mathcal{O}(100)$ Martin:2013tda ) for single field models of inflation. However, if we want to go beyond pure phenomenology to construct a particle-physics motivated models of inflation, one finds that the main restriction is that the potential must remain sufficiently flat for long enough for inflation to take place. In other words, not only should the classical potential have a very flat shape, it needs to be protected from quantum corrections. This is commonly referred to as the eta-problem in inflation: How do we keep quantum radiative corrections from generating a large mass term for the inflaton? Note that the theoretical obstruction comes from noting that the potential should be sensitive to Planck-suppressed operators, the absence of which would be an extreme fine-tuning problem. However, more importantly, current observations tell us that the primordial density perturbations have an adiabatic, Gaussian and near scale-invariant spectrum. This requires that the effective field theory (EFT) of inflation is that of a light scalar field, which is precisely what the above-mentioned eta problem makes it difficult to realize.
One way to achieve the above would be protecting the flatness of the potential via symmetries, such as supersymmetry (which would cancel the bosonic and fermionic contributions to the mass) or having an approximate shift symmetry (see, for instance, Baumann:2014nda ). The first suggestion works for highly fine-tuned superpotentials McAllister:2007bg , which just kicks the naturalness problem to a different starting point. Axion fields do posses weak shift symmetries Freese:1990rb , but these are ruled out from both theoretical (super-Planckian decay constants are ruled out by the Weak Gravity Conjecture Rudelius:2015xta ; Heidenreich:2015wga ) as well as from observational constraints Stein:2021uge . For inflationary models which require a large field excursion, the sensitivity of the potential to Planckian physics is enhanced, further worsening this problem. The broader point is that since there are no global symmetries in Quantum Gravity Banks:2010zn , this means all such models employing some symmetry would require a degree of fine-tuning to make them work. A recent, systematic manner of characterizing these problems is the so-called ‘Swampland program’ Palti:2019pca ; Bedroya:2019tba .
@@
The overarching problem of cold inflation is not that there does not exist a single model that still fits the bill as one can always reverse-engineer some potential, if required by adding more fields or additional symmetries or couplings, which can explain observed data (see Kallosh:2025rni ; Aoki:2025wld for examples post-ACT). The conceptual problem is how to screen inflation from the effects of trans-Planckian physics that come from the effects of pre-inflationary dynamics. One of the main qualities of inflation is that it is an attractor solution, and the standard Bunch-Davies initial conditions for the quantum fluctuations that source primordial inhomogeneities is a quantum attractor Kaloper:2018zgi for a quasi-de Sitter (dS) expansion. However, all of these nice qualities would be spoiled if one cannot justify how the inflaton mass remains small despite quantum corrections, as has been observed from the primordial power spectrum of density perturbations. And this requires a principled model, with a new symmetry that can be embedded from fundamental theory, which can explain why the potential remains approximately flat during inflation. Finally, note that even if such a mechanism is constructed, cold inflation would still need to invoke the reheating scenario via which the universe is repopulated giving rise to a radiation-dominated epoch at the end of inflation.
I.1 A new physical input: Dissipative corrections
@@
The above discussion implicitly ignores the interactions of the inflaton with other matter components during the quasi-dS expansion of the universe or, at best, assumes that these interactions only affect the potential through quantum corrections. However, one should also include dissipative that typically stem from interactions of the inflaton field. If the radiation production rate is non-negligible during inflation, this would compensate the super-cooling due to the exponential expansion and would not result in an empty universe. This is what has been shown to be the case in standard models of the warm inflation (WI) paradigm Berera:1995ie , which results in a radiation bath that is in quasi-thermal equilibrium. As there exists a radiation bath, the energy of the inflaton field can be dumped into this environment continuously through dissipative terms provided the microphysics causing dissipation has a characteristic time-scale that is much larger than the local expansion rate of the background spacetime Berera:1996nv ; Berera:1998gx . As is obvious from this description, there is no need for reheating in WI as it allows a smooth and gradual phase transition from exponential expansion to a radiation-dominated standard big-bang cosmology Berera:1996fm .
@@
Unlike in cold inflation, where the height of the potential determines the energy density of the universe while its slope fixes the slow-roll dynamics, there is an additional friction term in WI due to the radiation bath that keeps the universe warm despite the supercooling quasi-dS expansion. The background dynamics in WI is still dominated by the potential energy of the inflaton field, $V$. However, the inflaton field dynamics is not isolated and there exists the radiation energy density $\rho_{r}^{1/4}>H$, where $H$ is the approximately constant Hubble parameter during inflation. The resulting background dynamics in WI, for the inflaton and radiation energy density, modified by the presence of a dissipation term $\Upsilon$, are, respectively Berera:1995ie
$\ddot{\phi}+(3H+\Upsilon)\dot{\phi}+V_{,\phi}=0,$ (1)
$\dot{\rho}_{r}+4H\rho_{r}=\Upsilon\dot{\phi}^{2},$ (2)
@@
where $\Upsilon$ is the dissipation coefficient (for reviews on its derivation from first principles quantum field theory, see, e.g. Berera:2008ar ; Kamali:2023lzq and references there in) and it can be, in general, a function of both the temperature $T$ of the generated thermal radiation bath and on the inflaton amplitude $\phi$.
@@
$\epsilon_{V}$ $=$ $\frac{M_{\rm Pl}^{2}}{2}\left(\frac{V^{\prime}}{V}\right)^{2}<1+Q\,,$ (3)
$\eta_{V}$ $=$ $M_{\rm Pl}^{2}\left(\frac{V^{\prime\prime}}{V}\right)<1+Q\,$ (4)
@@
The presence of the dissipation term in WI not only modifies the background dynamics, but also strongly affects the perturbations, with in particular the inflaton and radiation perturbations becoming coupled due to the temperature dependence on $\Upsilon$ Berera:1995wh ; Berera:1999ws . In this case, the scalar of curvature of perturbations in WI is modified with respect to the CI case and it can be written as Ramos:2013nsa ; Benetti:2016jhf
$P_{\mathcal{R}}(k/k_{*})=\left(\frac{H^{2}}{2\pi\dot{\phi}}\right)^{2}\mathcal{F}(k/k_{*}),$ (5)
where all quantities are meant to be evaluated at Hubble radius crossing scale $k=aH$. In Eq. (5), $\mathcal{F}(k/k_{*})$ is defined as
$\mathcal{F}(k/k_{*})\equiv\left(1+2n_{*}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)G(Q),$
where $n_{*}$ accounts for the possibility of a thermal distribution of the inflaton field due to the presence of the radiation bath and assumes the Bose-Einstein distribution $n_{*}\equiv n_{\rm BE}=1/[\exp(H/T)-1]$ when fully thermalized (for intermediate cases, see, e.g. Ref. Bastero-Gil:2017yzb ). In Eq. (6), the function $G(Q)$ models the effect of the coupling of the inflaton and radiation fluctuations and which can only be determined numerically by solving the coupled system of perturbation equations in WI Graham:2009bf ; Bastero-Gil:2011rva ; Bastero-Gil:2014jsa . Equation (5) can be expanded in terms of the small scale $k$-dependence as
$P_{\mathcal{R}}(k/k_{*})\simeq A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1+\frac{\alpha_{s}}{2}\ln(k/k_{*})+\cdots},$ (7)
@@
and $\alpha_{s}$ is the running (the ellipses in Eq. (7) denote higher order running terms, which we neglect here),
$\alpha_{s}=\frac{dn_{s}(k/k_{*})}{d\ln(k/k_{*})}\Bigr{|}_{k\to k_{*}}.$ (9)
In WI, expressions for the tilt and running can be obtained explicitly by using Eq. (5) (see also, e.g.  Das:2022ubr ).
@@
The Planck+ACT+DESI data (P-ACT-LB hereafter) gives the best fit value of $n_{s}=0.974\pm 0.003$, which clearly shows that the spectrum is even more scale invariant. If one adds in the BICEP/Keck (P-ACT-LB-BK18 hereafter) data BICEP:2021xfz ; BICEPKeck:2024stm , one can turn this into a constraint on the second slow-roll parameter $\eta_{V}\sim\mathcal{O}(10^{-3})$ ACT:2025fju ; ACT:2025tim , which is very close to zero. Thus, the almost scale-invariance of the spectrum can be seen to make the eta-problem of cold inflation even worse, requiring extreme fine-tuning of the alpha-attractor models to include further arbitrary parameters of non-minimal couplings of the inflaton to gravity Kallosh:2025rni ; Kallosh:2013tua . On the other hand, this is not required in WI. For the strong dissipation regime, the value of $n_{s}$ is naturally led to larger values closer to one. The physical reason behind this is that a given wavelength mode attains its constant value slightly before it exits the horizon while the dissipative terms slows down the evolution of the field.
@@
Although quantum fluctuations are omnipresent in all models of inflation, Eq. (5) shows that the thermal and dissipation effects can dominate over them in WI and are primarily responsible for sourcing the adiabatic curvature perturbations. More interestingly, in the strong dissipative regime, where $Q>1$, these fluctuations freeze out even before they become superhorizon leading to an enhancement in the amplitude of the power spectrum. Since there is no such modification to the spectrum of the primordial tensor modes, this means that WI models would always lead to a suppression of the tensor-to-scalar ratio $r$ (noticing that dissipative effects to the gravitons are tiny for any temperature below the Planck scale). This implicitly is another reason why data seems to be favouring WI models, as the upper bound on $r$ is now constrained to be below $0.036$ BICEP:2021xfz . This will be further constrained in the near future by various missions such as LiteBird, BICEP/Keck, CMB-S4 etc. How large the value of $r$ we find would be the smoking gun in deciding between weak and strong dissipative regimes of WI, and would be crucial as an indication of whether the ‘natural’ models of WI, parametrized by a large value of $Q$, are preferred or not.
II It is time to ACT now

### v2

@@ Abstract
The recently released data from the Atacama Cosmology Telescope (ACT) confirms that the primordial scalar spectrum is extremely flat. This, together with current upper bounds on the tensor-to-scalar ratio, implies that the simplest models of inflation coming from particle physics (for instance, a minimally-coupled scalar with monomial potentials) need additional ingredients in order to make them compatible with observations. Instead of invoking arbitrary new couplings or new interactions that are not protected symmetries, we argue that dissipation of the inflaton field with the radiation bath should be added as a new physical principle. Accordingly, we show that warm inflation provides the correct paradigm to explain the current observations, given very natural choices of dissipative terms. The model analyzed here has mirror and $Z_{4}$ symmetries, which explicitly protect the inflaton potential from large quantum and thermal corrections. We use a recent precision numerical code designed for warm inflationary perturbations, improving on the determination of the cosmological observables previously obtained for such models.
1 Spoiled for choice, yet not a single ‘principled’ one
@@
The main prototype for inflation is a scalar field slowly rolling down an almost flat potential, such that the potential energy of the field dominates over its kinetic energy, and sources the Friedmann equation. The Hubble radius remains approximately constant during this phase, leading to an increase of the size of the universe by a factor of $e^{3N}$ if inflation lasted for NN $e$-folds. This results in a supercooled universe which is practically empty, and one needs a phase of reheating/preheating where the energy density of the inflaton is released by creating a shower of particles in radiation form.
The above story is simple enough that theorists have successfully built a zoo of potentials ($\sim\mathcal{O}(100)$ [1]) for single field models of inflation. However, if we want to go beyond pure phenomenology to construct a particle-physics motivated models of inflation, one finds that the main restriction is that the potential must remain sufficiently flat for long enough for inflation to take place. In other words, not only should the classical potential have a very flat shape, it needs to be protected from quantum corrections. This is commonly referred to as the eta-problem in inflation: How do we keep quantum radiative corrections from generating a large mass term for the inflaton? Note that the theoretical obstruction comes from noting that the potential should be sensitive to Planck-suppressed operators, the absence of which would be an extreme fine-tuning problem. However, more importantly, current observations tell us that the primordial density perturbations have an adiabatic, Gaussian and near scale-invariant spectrum. This requires that the effective field theory (EFT) of inflation is that of a light scalar field, which is precisely what the above-mentioned eta problem makes it difficult to realize.
One way to achieve the above would be protecting the flatness of the potential via symmetries, such as supersymmetry (which would cancel the bosonic and fermionic contributions to the mass) or having an approximate shift symmetry (see, for instance, [2]). The first suggestion works for highly fine-tuned superpotentials [3], which just kicks the naturalness problem to a different starting point. Axion fields do have weak shift symmetries [4], but these are ruled out from both theoretical (super-Planckian decay constants are ruled out by the Weak Gravity Conjecture [5, 6]) as well as from observational constraints [7]. For inflationary models which require a large field excursion, the sensitivity of the potential to Planckian physics is enhanced, further worsening this problem. The broader point is that since there are no global symmetries in Quantum Gravity [8], this means all such models employing some symmetry would require a degree of fine-tuning to make them work. A recent, systematic manner of characterizing these problems is the so-called ‘Swampland program’ [9, 10].
@@
This work presents several advancements over previous studies in multiple respects:
  1. 1.
    We begin by performing a new analysis of the warm little inflaton (WLI) model originally introduced in Ref. [28], employing the recently released WI2easy code [29]. This code enables a more efficient and precise computation of both the background dynamics and curvature perturbations, improving upon the accuracy of earlier numerical treatments. For the benchmark parameter set considered in Ref. [28], corresponding to a dissipation ratio $Q=\Upsilon/(3H)=0.27$, we obtain updated values for the spectral index and tensor-to-scalar ratio, $n_{s}=0.960$ and $r=9.5\times 10^{-4}$, respectively. These should be compared with the previous estimates $n_{s}=0.964$ and $r=8.0\times 10^{-4}$ reported in Ref. [28]. The variation in $n_{s}$ is already at the level of sensitivity of the precision data from ACT and Planck.
  2. 2.
    In addition, we re-examine, using the WI2easy framework, the results of Ref. [30], where dissipation arises from direct interactions between the inflaton and scalar fields, in contrast with the fermionic interactions originally considered in Ref. [28].
  3. 3.
    Beyond improving the precision in determining the power spectrum for warm inflation, we evaluate all quantities at the Hubble radius crossing point $N_{*}$ for the pivot scale. To this end, we follow the post-inflationary dynamics up to the onset of effective radiation domination, where the equation of state approaches $w=1/3$. This procedure allows for a more accurate estimation of observable quantities. In particular, while Refs. [28] and [30] presented results at a fixed number of e-folds, $N_{*}=60$, the more precise analysis performed here yields $N_{*}=58.27$ for the benchmark case studied in Ref. [28] ($Q_{*}=0.27$) and $N_{*}=55.86$ for the benchmark case studied in Ref. [30] ($Q_{*}=100$).
  4. 4.
    Furthermore, we extend the analysis to a unified WLI scenario that simultaneously incorporates both fermionic and scalar dissipative channels, performing a detailed study of the resulting dynamics and predictions. This is the first time that this extended and more complete model is studied.
  5. 5.
    All numerical analyses for the three models have been contrasted with the most recent ACT CMB data, as well as with the earlier Planck/BICEP measurements.
  6. 6.
    We also provide updated results for the running of the spectral index, which recent ACT observations suggest may take positive values. We examine this running separately for the fermionic, scalar, and combined interaction cases within the full WLI framework. Our results indicate that the recent CMB data from ACT show a clear preference for the strong dissipation regime of warm inflation.
  7. 7.
    Finally, we present, for the first time, a dedicated study of the possible thermalization of the inflaton in the extended WLI model, identifying parameter regions that favor such thermalization and discussing the resulting implications for the power spectrum and other observable quantities.
Throughout this paper, we employ the natural units: the speed of light, Planck’s constant, and Boltzmann’s constant are all set to $1$, $c=\hbar=k_{B}=1$, respectively. We consider a spatially flat homogeneous and isotropic background metric with scale factor $a(t)$, where $t$ is physical time. Dimensional quantities are expressed in terms of the reduced Planck mass, $M_{\rm Pl}=(8\pi G)^{-1/2}\simeq 2.44\times 10^{18}$ GeV, and where $G$ is Newton’s gravitational constant. The Hubble expansion rate is $H(t)={\dot{a}(t)}/a(t)$ and overdots mean derivative with respect to time.
The remainder of this paper is organized as follows. In section 2, we briefly review the warm inflation (WI) dynamics and the role of dissipation. In section 3, we explain the microphysics of the WI model used in our analysis. In section 4, we give all our numerical results and their connection to the recent ACT data. Our conclusions are given in section 5. Finally, in the appendix A we analyze the possibility of thermalization of the inflaton excitations in the model we have studied.
2 A new physical input: Dissipative corrections
@@
The above discussion implicitly ignores the interactions of the inflaton with other matter components during the quasi-dS expansion of the universe or, at best, assumes that these interactions only affect the potential through quantum corrections. However, one should also include dissipative effects that typically stem from interactions of the inflaton field. If the radiation production rate is non-negligible during inflation, this would compensate the super-cooling due to the exponential expansion and would not result in an empty universe. This is what has been shown to be the case in standard models of the WI paradigm [34], which results in a radiation bath that is in quasi-thermal equilibrium. As there exists a radiation bath, the energy of the inflaton field can be dumped into this environment continuously through dissipative terms provided the microphysics causing dissipation has a characteristic time-scale that is much larger than the local expansion rate of the background spacetime [35, 36]. As is obvious from this description, there is no need for reheating in WI as it allows a smooth and gradual phase transition from exponential expansion to a radiation-dominated standard big-bang cosmology [37].
@@
Unlike in cold inflation, where the height of the potential determines the energy density of the universe while its slope fixes the slow-roll dynamics, there is an additional friction term in WI due to the radiation bath that keeps the universe warm despite the supercooling quasi-dS expansion. The background dynamics in WI is still dominated by the potential energy of the inflaton field, $V$. However, the inflaton field dynamics is not isolated and there exists the radiation energy density $\rho_{r}^{1/4}>H$, where $H$ is the approximately constant Hubble parameter during inflation. The resulting background dynamics in WI, for the inflaton and radiation energy density, modified by the presence of a dissipation term $\Upsilon$, are, respectively [34]
$\ddot{\phi}+(3H+\Upsilon)\dot{\phi}+V_{,\phi}=0,$ (2.1)
$\dot{\rho}_{r}+4H\rho_{r}=\Upsilon\dot{\phi}^{2},$ (2.2)
@@
where $\Upsilon$ is the dissipation coefficient (for reviews on its derivation from first principles quantum field theory, see, e.g. [38, 39] and references there in) and it can be, in general, a function of both the temperature $T$ of the generated thermal radiation bath and on the inflaton amplitude $\phi$.
@@
$\epsilon_{V}$ $=$ $\frac{M_{\rm Pl}^{2}}{2}\left(\frac{V^{\prime}}{V}\right)^{2}<1+Q\,$, (2.3)
$\eta_{V}$ $=$ $M_{\rm Pl}^{2}\left(\frac{V^{\prime\prime}}{V}\right)<1+Q,$ (2.4)
@@
The presence of the dissipation term in WI not only modifies the background dynamics, but also strongly affects the perturbations, with in particular the inflaton and radiation perturbations becoming coupled due to the temperature dependence on $\Upsilon$. The scalar primordial perturbations for warm inflation in the thermally induced case were obtained here [40, 41]. Thus, the scalar perturbation power spectrum in WI is modified with respect to the CI case and it can be written as [42, 43]
$P_{\mathcal{R}}(k/k_{*})=\left(\frac{H^{2}}{2\pi\dot{\phi}}\right)^{2}\mathcal{F}(k/k_{*}),$ (2.5)
where all quantities are meant to be evaluated at Hubble radius crossing scale $k=aH$. In eq. (2.5), $\mathcal{F}(k/k_{*})$ is defined as
$\mathcal{F}(k/k_{*})\equiv\left(1+2n_{*}+\frac{2\sqrt{3}\pi Q}{\sqrt{3+4\pi Q}}\frac{T}{H}\right)G(Q).$
Here the last term in the parenthesis arises from the thermalized heat bath inducing fluctuations of the inflation field through the fluctuation-dissipation dynamics. In addition, the inflaton field can also become thermalised and $n_{*}$ accounts for the possibility of a thermal distribution of the inflaton field with a Bose-Einstein distribution $n_{*}\equiv n_{\rm BE}=1/[\exp(H/T)-1]$ when fully thermalized (for intermediate cases, see, e.g. ref. [44, 45]). In the appendix A, we discuss the possibility of thermalization of the inflaton and give a justification of when it is appropriate to consider $n_{*}\equiv n_{\rm BE}$ in eq. (2.6). Finally the first term in the above parenthesis, the $1$, reflects the contribution of the ground state vacuum fluctuations. In eq. (2.6), the function $G(Q)$ models the effect of the coupling of the inflaton and radiation fluctuations and which can only be determined numerically by solving the coupled system of perturbation equations in WI [46, 47, 48] (see also refs. [49, 29] for public codes that are dedicated to evolve the full set of perturbation equations in WI and to numerically determine $G(Q)$ for different forms of dissipation coefficients). Equation (2.5) can be expanded in terms of the small-scale $k$-dependence as
$P_{\mathcal{R}}(k/k_{*})\simeq A_{s}\left(\frac{k}{k_{*}}\right)^{n_{s}-1+\frac{\alpha_{s}}{2}\ln(k/k_{*})+\cdots},$ (2.7)
@@
and $\alpha_{s}$ is the running (the ellipses in eq. (2.7) denote higher order running terms, which we neglect here),
$\alpha_{s}=\frac{dn_{s}(k/k_{*})}{d\ln(k/k_{*})}\Bigr|_{k\to k_{*}}.$ (2.9)
In WI, expressions for the tilt and running can be obtained explicitly by using eq. (2.5) (see also, e.g.  [50]).
@@
The Planck+ACT+DESI data (P-ACT-LB hereafter) gives the best fit value of $n_{s}=0.974\pm 0.003$, which clearly shows that the spectrum is even more scale invariant. If one adds in the BICEP/Keck (P-ACT-LB-BK18 hereafter) data [51, 52], one can turn this into a constraint on the second slow-roll parameter $\eta_{V}\sim\mathcal{O}(10^{-3})$ [23, 22], which is very close to zero. Thus, the almost scale-invariance of the spectrum can be seen to make the eta-problem of cold inflation even worse, requiring extreme fine-tuning of the alpha-attractor models to include further arbitrary parameters of non-minimal couplings of the inflaton to gravity [31, 53]. On the other hand, this is not required in WI. For the strong dissipation regime, the value of $n_{s}$ is naturally led to larger values closer to one. The physical reason behind this is that a given wavelength mode attains its constant value slightly before it exits the horizon, whereas the dissipative terms slow down the evolution of the field. In addition to slightly larger values for the scalar tilt, the recent ACT data also indicates a greater preference for a positive running, with $\alpha_{s}=0.0062\pm 0.0052$, while the Planck results pointed toward a sllightly negative value [25, 24], $\alpha_{s}=-0.0041\pm 0.0067$. As we shall see, warm inflation can easily accommodate both of these results, in particular in the so-called strong dissipative regime $Q>1$.
@@
There is both a conceptual and practical importance of having warm inflation in the strong dissipative regime. Although quantum fluctuations are omnipresent in all models of inflation, eq. (2.5) shows that the thermal and dissipation effects can dominate over them in WI and are primarily responsible for the sourcing adiabatic curvature perturbations. More interestingly, in the strong dissipative regime, where $Q>1$, these fluctuations freeze out even before they become superhorizon leading to an enhancement in the amplitude of the power spectrum. Since there is no such modification to the spectrum of the primordial tensor modes, this means that WI models would always lead to a suppression of the tensor-to-scalar ratio $r$ (noticing that dissipative effects to the gravitons are tiny for any temperature below the Planck scale). This implicitly is another reason why the data appear to favor WI models, as the upper bound on $r$ is now constrained to be below $0.036$ [51]. This will be further constrained in the near future by various missions such as LiteBird, BICEP/Keck, CMB-S4 etc. How large the value of $r$ we find would be the smoking gun in deciding between weak and strong dissipative regimes of WI, and would be crucial as an indication of whether the ‘natural’ models of WI, parametrized by a large value of $Q$, are preferred or not.
3 It is time to ACT now